User Tools

Site Tools


lbaops:lbamar2007:vt11c3holog

eVLBI experiment

The log from experiment vt11c3 has been copied to: ftp.atnf.csiro.au:/pub/people/vlbi/incoming/

ANTAB and UVFLG files have been copied to the same place.

Additional experiment notes:

Missing PPS Errors fixed prior to this experient by rerouting DAS to VSIB cable in an individual screen and running the cable along the floor instead of in the duct that goes under the floor and contains lots of power cables.

2007.083.18:28:39.15;“DAS VSOP_HO

2007.083.18:29:22.55;”Agilent 4.2 GHz

2007.083.18:29:47.54;“SMY 502

2007.083.18:30:14.25;”coherence 1639

2007.083.18:31:27.63;“DAS coonnected to VSIC via a newly screened cable.

2007.083.18:32:19.60;”observer Aidan and Claire Hotan, Brett Reid

2007.083.18:36:07.15;“We are ready for eVLBI

2007.083.19:36:15.28;”switch shows data is being extracted. hopefully at a huge rate

2007.083.19:36:39.16;“weather is calm, clear, cold

2007.083.19:36:54.77;”sun starting to rise

regards,

Brett.

lbaops/lbamar2007/vt11c3holog.txt · Last modified: 2015/12/18 16:38 (external edit)